OpenStudy (summersnow8):

(1) A bullet with mass 50.0 g is fired with an initial velocity of 625 m/s from a gun with a mass 3.50 kg. What is the speed of recoil (in m/s) of the gun? (2) How many Newtons of force are required to stop a truck 2250 kg truck going 32.0 km/h in 4.00s?

OpenStudy (shamim):

1) Use conservation of momentum law

OpenStudy (shamim):

m1*u1+m2*u2=m1*v1+m2*v2

OpenStudy (summersnow8):

what does u stand for?

OpenStudy (summersnow8):

could you work me through it? i'm not too familiar with that formula

OpenStudy (shamim):

m1=mass off bullet=50g=?kg m2=mass of gun=3.5kg u1=initial velocity of bullet=0 u2=initial velocity of gun=0 v1=final velocity of bullet =625m/s v2=?

OpenStudy (shamim):

v2=recoil speed of gun

OpenStudy (shamim):

No

OpenStudy (summersnow8):

(0.05 kg)(625 m/s)+(3.5 kg)(0 m/s)= (0.065 kg)(625 m/s)+ (3.5 kg)v2

OpenStudy (shamim):

m1=50g=50/1000kg=.05kg Right?

OpenStudy (rajat97):

i think it should be (0.05 kg)(0 m/s)+(3.5 kg)(0 m/s)= (0.06 kg)(625 m/s)+ (3.5 kg)v2

OpenStudy (shamim):

.05*0+3.5*0=.05*625+.05*v2 v2=?

OpenStudy (shamim):

Why .06 kg?

OpenStudy (rajat97):

oh! i'm sorry

OpenStudy (rajat97):

*0.05kg

OpenStudy (summersnow8):

i got 31.25 = 31.25 + 3.5 V2

OpenStudy (rajat97):

no you may be using the wrong equation the correct one is 05*0+3.5*0=.05*625+.05*v2 as @shamim said

OpenStudy (shamim):

.05*0=0 Right?

OpenStudy (rajat97):

so you should get -31.25=3.5v2 and i posted the wrong equation it should be 05*625+3.5*v2

OpenStudy (shamim):

3.5*9=0 Right?

OpenStudy (shamim):

0+0=0 Right?

OpenStudy (summersnow8):

V2 = -8.9285 m/s, but since speed is scalar it is just 8.9285

OpenStudy (shamim):

Left hand side of my equation will b 0 Right?

OpenStudy (shamim):

- reveals v1 nd v2 r oposit in direction

OpenStudy (shamim):

Ur bullet is going to forward nd ur gun is going to backward

OpenStudy (summersnow8):

how do I solve #2 now?

OpenStudy (shamim):

Who will get medal?????!!!!!

OpenStudy (summersnow8):

whoever helps me solve both problems

OpenStudy (summersnow8):

How many Newtons of force are required to stop a truck 2250 kg truck going 32.0 km/h in 4.00s?

OpenStudy (shamim):

U should post ur second question again in another post

OpenStudy (summersnow8):

ok

OpenStudy (rajat97):

thanks for the medal @Summersnow8

OpenStudy (summersnow8):

@rajat97 no problem